in 1990, the life expectancy of males in a certain country was 63.1 years.in 1996, it was 65.4 years.let E represent the life expectancy in year t and let t represent the number of years since 1990.
the linear function E(t) that fits the data is E(t)=?
use the function to predict the life expectancy in males in 2007
E(17)=? all answers must be rounded to the nearest tenth

Answer by oscargut(686)

E(t)=at+b, we have to find a and b
in 1990, the life expectancy of males in a certain country was 63.1
then E(0)=63.1 then E(0)=a*0+b=b=63.1
then E(t)=at+63.1
in 1996, it was 65.4 years
then E(6)=65.4 then 6a+63.1=65.4
then 6a=2.3
a= 2.3/6
then E(t)=(2.3/6)t+63.1
the life expectancy in males in 2007 is E(7)=(2.3/6)*7+63.1=65.783333=65.8(aprox)
E(17)=(2.3/6)17+63.1=128.26666=128.3 (aprox)